A pharmacist wishes to mix a solution that is %8 Minoxidil. She has on hand 90 ml of a 1% solution and wishes to add some 10% solution to obtain the desired 8% solution. How much 10% solution should she add?
Answer by stanbon(75887)

A pharmacist wishes to mix a solution that is %8 Minoxidil. She has on hand 90 ml of a 1% solution and wishes to add some 10% solution to obtain the desired 8% solution. How much 10% solution should she add?
----------------------------
Equation:
mino + mino = mino
0.01*90 + 0.10(x) = 0.08(90+x)
---
Multiply thru by 100 to get:
1*90 + 10x = 8*90 + 8x
---
2x = 7*90
x = 7*45 = 315 ml (amt of 10% solution needed)
===============================
Cheers,
Stan H.